Starting from rest, a car accelerates at 2.4 m/s2 up a hill that is inclined 5.5° above the horizontal. How far horizontally and vertically has the car traveled in 9 s
Answer #1
Horizontally:

x = 1/2*a*t^2

x = 1/2*2.4*(cos 5.5)*9^2

Vertically

y = 1/2*a*t^2

y = 1/2*2.4*(sin 5.5)*9^2

PLEASE RATE.